# Modeling the amplitude and energy decay of a weakly damped harmonic oscillator using the energy dissipation rate and a simple trick

### v1

@@ V A simpler version of our model
Exponential amplitude decay and purely exponential approximation of the energy decay of weakly damped harmonic oscillator have recently been derived using the energy dissipation rate averaged over one period of the corresponding undamped system, i.e. $T_{0}$, with the approximation that the amplitude remains constant over time intervals $\Delta t\sim T_{0}$ [8]. In our model, this approximation corresponds to neglecting terms with $\dot{f}(t)$ in the velocities (6) and (7), while the displacements (4) and (5) remain the same. In this case, it is easy to show that the modeled energies for both pairs of initial conditions are the same, i.e.
@@
and the corresponding dissipation rates are
@@
Now, it is even easier to see that the sum of (25) and (26) leads to equation (14). Thus, we obtain $f(t)=e^{-\frac{b}{2m}t}$ again, but in this case, we only get the exponential energy decay, i.e., $E_{1}(t)=E_{2}(t)=E_{0}e^{-\frac{b}{m}t}$, which is a good approximation of the exact energies (23), with $\frac{b}{2m}\ll\omega_{0}$, only if one is interested in behavior of the energy on time scales greater than $T_{0}$. This simplified version of our derivation is suitable for high school students, i.e., from a mathematical point of view, it is at a slightly lower level of difficulty than the method presented in [8] since there is no need to calculate the time integral of the energy dissipation rate if our trick is used.
VI Conclusion
@@
We showed how to obtain an excellent approximation of the solutions and the energy of a weakly damped harmonic oscillator without solving the differential equation of the second order, i.e., equation (1), and without any previous insights into the analytical form of the solutions of that equation. The model we introduced and the trick we used are suitable for first-year undergraduates. In addition, we commented on a simplified version of our model and derivation, which are more suitable for high school students.
VII Acknowledgments

### v2

@@ V A simpler approach
Exponential amplitude decay and purely exponential approximation of the energy decay of weakly damped harmonic oscillator have recently been derived, in a manner suitable for high school students, using the energy dissipation rate averaged over one period of the corresponding undamped system, i.e. $T_{0}$, with the approximation that the amplitude remains constant over time intervals $\Delta t\sim T_{0}$ Kontomaris2024 . In our model, this approximation corresponds to neglecting terms with $\dot{f}(t)$ in the velocities (6) and (7), while the displacements (4) and (5) remain the same. In this case, it is easy to show that the modeled energies for both pairs of initial conditions are the same, i.e.
@@
and the corresponding dissipation rates (3) are
@@
Now, it is even easier to see that the sum of (26) and (27) leads to equation (14). Thus, we obtain $f(t)=e^{-\frac{b}{2m}t}$ again, but in this case, we only get the exponential energy decay, i.e., $E_{1}(t)=E_{2}(t)=E_{0}e^{-\frac{b}{m}t}$, which is a good approximation of the exact energies (23), with $\frac{b}{2m}\ll\omega_{0}$, only if one is interested in the behavior of the energy on time scales greater than $T_{0}$. This simplified version of our derivation is suitable for high school students, i.e., from a mathematical point of view, it is at a slightly lower level of difficulty than the method presented in Kontomaris2024 since there is no need to calculate the time integral of the energy dissipation rate if our trick is used.
VI Incorporating our approach to teaching
As we mentioned in the introduction, there are two typical approaches to this topic in standard physics textbooks for undergraduate studies:
  • The general exact solution of differential equation (1) is just given Resnick10 ; Young2020university , without derivation, along with the expression for damped angular frequency. We will refer to this approach as approach A.
  • A solution of the form $x(t)=Ae^{-\frac{t}{2\tau}}\cos(\omega t+\phi)$ is assumed Waves , without explaining why the exponential function is taken and why the factor $1/(2\tau)$ in the exponent, then, by direct substitution in (1), $\omega$ and $\tau$ are determined for which the assumed solution satisfies (1). We will refer to this approach as approach B.
Here, we comment on the benefits of having the instructor first introduce students to our modeling of the solutions and energies of a weakly damped harmonic oscillator and then continue with approach A or approach B.
Through section III, students can learn that the amplitude decays exponentially in the weak damping regime and that the relationship between $\frac{b}{2m}$ and $\omega_{0}$, i.e. between $b$ and $2\sqrt{km}$, characterizes the damping strength, and not the magnitude of $b$ per se. These findings are not given to them as finished products without derivation and without explanation, i.e., students can adopt them with understanding because our approach uses a mathematical formalism with which they are familiar and comprehensible approximations. Furthermore, instructors can present figures with our modeled energies, as Fig. 4, without comparing them to exact energies. By considering such figures, students can connect plateaus in modeled energies with the turning points, and the importance of the monotonous decrease of energy can be pointed out to them, i.e., the energy of free-damped oscillations cannot increase. This last point is sometimes overlooked even by experienced authors, e.g., in Kontomaris2024 ; Dourmashkin shown examples of exact energies of free damped harmonic oscillator display visible local maxima, which immediately indicates some plotting error, and can be potentially confusing to students. Before continuing with approach A or B, an additional helpful step would be to show a figure such as Fig. 3. It would be sufficient to show the behavior for initial conditions with a purely potential initial energy for students to see that, for larger values of $\frac{b}{2m\omega_{0}}$, both frequency and phase begin to deviate from the frequency and phase of the undamped case. That the phase of the damped solution must be changed compared to the phase of the undamped solution is also evident from the discussion in section III.
If the instructor now continues with approach A, i.e., gives students the general exact solution of (1) and the expression for damped angular frequency, even without derivation, students will have a clear understanding of the exponential function in the exact solution and intuition about the frequency decrease of damped oscillations in comparison to the undamped case. If the instructor continues with the approach B, he (or she) can try a solution of the form
$x(t)=Ae^{-\frac{b}{2m}t}\cos(\omega t+\phi)\,,$ (28)
and by direct substitution in (1) find that damped angular frequency has to be $\omega=\sqrt{\omega_{0}^{2}-\left(\frac{b}{2m}\right)^{2}}$, i.e. students will have clear idea why the instructors takes this form of the solution, and there is no need for introducing unknown and unintuitive factor $1/(2\tau)$ in the exponential function. At his point in both scenarios, the instructor can define the damping coefficient $\gamma=\frac{b}{2m}$, for simplicity, and specify that the damped system oscillates only if $\gamma<\omega_{0}$, i.e. in the underdamped regime, and, depending on the course program, further comment on the critical damping, i.e. $\gamma=\omega_{0}$, and overdamping, i.e. $\gamma>\omega_{0}$, regimes.
We note here that after the students are familiar with the solution (28), the instructors can obtain the exact solutions of the critical and overdamped regimes without ever formally solving the differential equation (1), i.e. the solutions of those two regimes are already contained in solution (28). To see that, one proceeds as follows. Solution (28) can be written as
$x(t)=e^{-\gamma t}\left(B\cos(\omega t)+C\sin(\omega t)\right)\,,$
and when we determine the constants $B$ and $C$, from initial conditions $x_{0}$ and $v_{0}$, we get
$x(t)=e^{-\gamma t}\left(x_{0}\cos(\omega t)+\frac{v_{0}+\gamma x_{0}}{\omega}\sin(\omega t)\right)\,.$ (29)
We can easily obtain the solution of critical damping regime by taking the $\gamma\rightarrow\omega_{0}$ limit of (29), since in that case $\omega\rightarrow 0$, $\lim_{\omega\rightarrow 0}\cos(\omega t)=1$ and $\lim_{\omega\rightarrow 0}\omega^{-1}\sin(\omega t)=t$. To obtain the solution of the overdamped regime, one can write $\omega=i\Omega$, where $\Omega=\sqrt{\gamma^{2}-\omega_{0}^{2}}$ and $i=\sqrt{-1}$, and use the connection of trigonometric and hyperbolic functions, i.e. $\sin(i\Omega t)=i\sinh(\Omega t)$ and $\cos(i\Omega t)=\cosh(\Omega t)$, in (29), as was done in Waves .
VII A short comment on experimental backup for our approach
Fig. 1 is one of the central ingredients of our approach. It gives us insight into the approximations we can use for weak damping. If we wanted to obtain the data for such a figure experimentally, the only tricky part would be to realize two different pairs of initial conditions with the same initial energy, one with purely potential initial energy and the other with purely kinetic initial energy. Here, we comment on only one simple idea to achieve a good approximation of such initial conditions, which even students without prior knowledge of damped oscillators could understand and apply in the laboratory. As a concrete example, we will consider a physical pendulum with linear damping due to eddy currents Wang ; Gonzalez2006 .
In such systems, all damping regimes can be achieved, from weak damping with $\gamma\sim 0.01\omega_{0}$ Wang to overdamped regime with $\gamma\gtrsim\omega_{0}$ Gonzalez2006 . Let us imagine that we are considering such a pendulum in the regime of small oscillations and weak damping with $\gamma\sim 0.01\omega_{0}$, which corresponds to the experimental conditions in Wang . We displace the pendulum by the angle $\theta_{0}$, let it swing freely for some time and record the data. This way, we would obtain data similar to the one shown as a red curve in Fig. 1(a), with angle as a displacement. Then, by comparison with the undamped case with the same initial conditions, we could easily see that the experimental data and undamped solution mostly overlap for $20\leq t\leq T_{0}/$ 2, with only slightly smaller amplitude of damped displacement at the first turning point, i.e. at $t=T_{0}/2$. Thus, we could conclude that the energy loss due to damping is negligible in the first quarter of a period, i.e. for $40\leq t\leq T_{0}/$ 4. If we then repeat the experiment with initial displacement $-\theta_{0}$ and plot the angle vs. time data starting from $t=T_{0}/4$, we would get a curve similar to a blue curve in Fig. 1(b), i.e., a curve which corresponds to initial conditions with purely kinetic initial energy which is negligibly smaller than the initial potential energy in the previous experiment.
VIII Conclusion
@@
We showed how to obtain an excellent approximation of the solutions and the energy of a weakly damped harmonic oscillator without solving the differential equation of the second order, i.e., equation (1), and without any previous insights into the analytical form of the solutions of that equation. The model and trick we introduced are suitable for first-year undergraduates. We commented on a simplified version of our model and derivation, which are more appropriate for high school students. In addition, we gave suggestions on integrating our approach, method, and results into teaching. We commented on obtaining the exact solutions for all damping regimes without using formal methods for solving second-order differential equations.
IX Acknowledgments